# ImFace++: A Sophisticated Nonlinear 3D Morphable Face Model with Implicit Neural Representations

### v2

@@ Abstract
Accurate representations of 3D faces are of paramount importance in various computer vision and graphics applications. However, the challenges persist due to the limitations imposed by data discretization and model linearity, which hinder the precise capture of identity and expression clues in current studies. This paper presents a novel 3D morphable face model, named ImFace++, to learn a sophisticated and continuous space with implicit neural representations. ImFace++ first constructs two explicitly disentangled deformation fields to model complex shapes associated with identities and expressions, respectively, which simultaneously facilitate the automatic learning of correspondences across diverse facial shapes. To capture more sophisticated facial details, a refinement displacement field within the template space is further incorporated, enabling a fine-grained learning of individual-specific facial details. Furthermore, a Neural Blend-Field is designed to reinforce the representation capabilities through adaptive blending of an array of local fields. In addition to ImFace++, we have devised an improved learning strategy to extend expression embeddings, allowing for a broader range of expression variations. Comprehensive qualitative and quantitative evaluations demonstrate that ImFace++ significantly advances the state-of-the-art in terms of both face reconstruction fidelity and correspondence accuracy.
@@ 1 Introduction
In 3DMMs, the most fundamental issue lies in the way to generate latent morphable representations, and during the past two decades, along with data improvement in scale, diversity, and quality [8, 11, 6, 41], remarkable progresses have been achieved. The methods are initially linear model based [1, 57, 58] and further extended to multilinear model based [9, 7, 56], where different modes are individually encoded. Unfortunately, for the relatively limited representation ability of linear models, these methods are not so competent at handling the cases with complicated variations, e.g. , exaggerated expressions. In the era of deep learning, a number of nonlinear models have been investigated with the input of 2D images [12, 19] or 3D meshes [14, 15, 16, 17, 44] by using Convolutional Neural Networks or Graph Neural Networks. They indeed deliver performance gains; however, restricted by the resolution of discrete representing strategies on input data, facial priors are not sufficiently captured, incurring loss of shape details. Besides, all current methods are dependent on the preposed procedure of point-to-point correspondence [10, 59, 60, 43], but face registration itself remains challenging.
@@
Figure 1: ImFace++ effectively encodes the intricate face variations associated with identities and expressions by two explicitly disentangled deformation fields with respect to a template face. Leveraging the established correspondence, it applies an additional refinement displacement field to capture high-frequency details in the individual-independent template space. This approach yields sophisticated and implicit representations that are highly adaptable for 3D face modeling.
@@
Recently, several studies have illuminated the potential of Implicit Neural Representations (INRs) [20, 21, 22, 26, 27] in the precise modeling of 3D geometries. INRs describe input observations as low-dimensional shape embeddings and estimate the Signed Distance Function (SDF) or the occupancy values of query points, allowing for the definition of surfaces with arbitrary resolution and topology through isocontours. Such continuous representations offer advantages over the traditional discrete ones like voxels, point clouds, and meshes, and have reported substantial success in shape reconstruction [62, 61, 28, 63], as well as surface registration [23, 25, 24]. Such an advantage suggests an alternative to 3DMM that can fulfill accurate correspondence and geometry deformation modeling in a unified network. Despite that, applying INRs to face modeling presents unique challenges. Unlike the objects with distinct shapes and limited non-rigid variations, such as indoor scenes and human bodies, facial surfaces are inherently similar but also encompass intricate deformations arising from the interweaving of multiple identities and rich expressions. Current INR methods struggle to precisely capture and disentangle these intricacies, as evidenced by the preliminary attempt  [33].
@@
On the other side, it is essential to emphasize the faithful reproduction of facial details. In the context of geometry modeling, there has been a growing interest in hybrid approaches that combine explicit and implicit representations to harness the strengths of both techniques [74, 72, 69], aiming to address the spectral bias [65] issue often observed in neural networks and learn more sophisticated geometry clues. However, incorporating explicit 3D representations into INRs can introduce discontinuities, potentially leading to training instability and over-fitting [91], especially when applied to complex structures like facial surfaces. An alternative research trend focuses on incorporating computer graphics techniques, such as displacement mapping [82, 83], into INRs to encode rich details. While these methods have shown remarkable performance in reconstructing single static scenes, the challenge remains in developing INR models capable of effectively capturing dynamically changing high-frequency details across multiple facial shapes, such as wrinkles and dimples.
@@
In response to the aforementioned challenges, this paper presents a novel two-stage 3D face morphable model, namely ImFace++, which substantially upgrades conventional 3DMMs by integrating customized INRs. In its initial stage, ImFace++ tackles nonlinear facial deformation modeling and establishment of correspondence among different individuals by creating separate INR sub-networks, which explicitly disentangle shape morphs into two deformation fields for identity and expression, respectively (as Fig. 1 shows). The template shape is defined by a deep SDF, therefore eliminating the need for the face registration phase required by numerical solutions. To build more sophisticated 3DMMs, ImFace++ further incorporates a refinement displacement field (RDF) in the second stage to capture high-frequency geometry details present in various faces. Notably, this stage for detail modeling is specifically tailored to adapt to the dynamic nature of facial expressions. By leveraging the established correspondences, RDF is learned within the individual-independent template space characterized by its relative stability, allowing the model to concentrate on encoding subtle details. Overall, by learning the disentangled deformation fields and RDF, ImFace++ accurately represents inter-individual changes and fine-grained clues while retaining the flexibility for its application in related tasks.
@@
Inspired by linear blend skinning [37], we have further introduced a Neural Blend-Field to decompose the entire facial deformation or geometry into semantically meaningful regions encoded by an array of local implicit functions and adaptively blend them through a lightweight module, leading to more advanced representations with reduced parameter overhead. In addition to ImFace++, an improved embedding learning strategy is devised to extend the latent space of expressions to allow a more diverse expression range.
More concisely, this study makes the following contributions:
@@
  • A novel INR-based 3DMM, namely ImFace++, which encodes complex facial shape variations associated with identity and expression through explicitly disentangled deformation fields. It enables the learning of flexible and semantically meaningful representations.
@@
  • A refinement displacement field (RDF) within the individual-independent template space. Leveraging the established correspondence, this field faithfully encodes high-frequency details, thereby enhancing the representation capabilities of ImFace++.
@@
  • A Neural Blend-Field, which enables fine-grained learning by adaptively blending an array of local implicit functions.
@@
  • Extensive experiments evidence that ImFace++ excels in synthesizing high-quality 3D faces with intricate details, outperforming the state-of-the-art alternatives.
@@
A preliminary version of this study was ImFace, previously published in [87]. This paper significantly improves it in the following aspects: (i) We extend the model to a two-stage framework. To sufficiently capture facial priors, the upgraded model effectively leverages the correspondences between different individuals, achieved by employing disentangled deformation fields in the first stage as in ImFace [87]; it further investigates a newly designed refinement field to progressively model face geometry, significantly alleviating the loss of shape details while relieving the burden of model learning. (ii) We update the model to reconstruct and synthesize 3D faces spanning a wide range of facial expressions and it renders more vivid results. We have also refined evaluations in both the observed and latent representation spaces, along with more comprehensive discussions. (iii) To further validate the proposed approach, we conduct comparisons with other INR-based methods [33, 66] in terms of 3D face reconstruction and face correspondence. (iv) We showcase applications in various scenarios, including facial expression editing and detail transfer, to demonstrate the adaptability of ImFace++ in the realm of 3D face modeling.
@@
The rest of this paper is organized as follows. Section 2 reviews related work on 3DMMs and INRs. Section 3 elaborates on the proposed ImFace++ method. Section 4 presents and analyzes the experimental results using the FaceScape database, followed by Section 5, which concludes this paper and provides insights for future perspectives.
@@ 2 Related Work
The published studies closely related to our study can be primarily summarized into: (i) 3D Morphable Models, (ii) Implicit Neural Representations, offering innovative methods for capturing complex 3D structures, and (iii) Fine-Grained Neural Representations for Geometry Modeling, delving into studies that specifically address the creation and application of fine-grained neural representations for geometry modeling. In the following, we provide a concise review of the relevant literature.
@@
The concept of 3D Morphable Face Models (3DMMs) was first introduced by Blanz and Vetter [1] as a foundational framework for facial representation. This approach involved registering a known template mesh to all training scans using the Non-rigid Iterative Closest Point (NICP) algorithm [40]. Principal Components Analysis (PCA) was subsequently employed to create prior face distributions. To account for identity-dependent expressions, 3DMMs were extended to multilinear models [9, 7, 56]. Over time, significant progress has been made aided by improvements in data quality [8, 11, 6, 41]. Notably, the Facial LAndmarks and More Expressions (FLAME) model [6] was introduced as an expressive model, capable of controlling facial expressions through a combination of jaw articulation and linear expression blendshapes. Trained on a comprehensive 3D facial dataset, including D3DFACS [42], FLAME delivered remarkable results, although it faced challenges in accurately capturing nonlinear facial deformations.
The development of deep neural networks has significantly accelerated the development of more potent nonlinear 3DMMs. A number of models have been trained directly from 2D images [12, 19, 18], but they often struggled to capture high-fidelity and fine-grained details. This limitation is attributed to the inherent challenges posed by the low resolution of input images in solving this ill-posed inverse problem. Recognizing the need to better leverage 3D face scans, Bagautdinov and colleagues [14] devised a method for mapping 3D meshes to the 2D space. Subsequently, several other studies [44, 15, 17, 16] directly pursued the learning of 3DMMs from meshes, employing spectral or spiral convolutions. While promising, these neural networks were established on discrete 3D representations, which posed limitations when confronted with intricate facial deformations.
@@ 2.2 Implicit Neural Representations
In recent years, implicit neural functions have emerged as powerful and effective representations for 3D geometry [20, 21, 22, 26, 27, 45]. These approaches have revolutionized the field by enabling continuous and highly expressive representations of geometric structures without the need for discretization. To address the complexities of shape variations and correspondence relationships, some methods have further incorporated an additional implicit deformation latent space [23, 24].
A number of studies have focused on applying INRs to model human heads or bodies [35, 34, 33, 32, 31, 36, 30, 47, 66, 88, 89, 90, 92, 93, 95, 94]. In particular, NPMs [95] and GNPM [94] have introduced approaches employing two neural networks to separately model body shape and pose spaces. Models such as H3D-Net [34] and SIRA++[92] have learned implicit priors for 3D head reconstruction. i3DMM[33] made significant strides by pioneering a generative implicit 3D morphable model tailored for human heads. However, utilizing neural networks to generate fine-grained details from raw signals has often resulted in over-smoothed or implausible facial shapes in these studies. A concurrent work, NPHM [66], has devised a parametric head model using local implicit fields, similar to our approach. It employed local INRs for identity representation and a globally conditioned INR for modeling expression deformations. Nonetheless, its ability for capturing high-frequency facial details was constrained by the globally conditioned neural network architecture it employed. Another series of works [88, 89, 90, 93] have concentrated on modeling facial appearance, leveraging techniques like differential volume rendering [49] and auto-decoding [20] to learn morphable radiance fields. However, these models represented the underlying geometry using density, leading to noisy geometry approximations.
Figure 2: ImFace++ overview. The model is constructed through two-stage learning. (a) The first stage primarily focuses on nonlinear facial deformation modeling and correspondence establishment between different faces. The network comprises three Mini-Nets blocks designed to explicitly disentangle shape morphs into separate deformation fields, where the Expression and Identity Mini-Nets blocks are associated with expression and identity deformations, respectively. Meanwhile, the Template Mini-Nets block learns the SDF of a template face shape. (b) With the correspondence established, ImFace++ proceeds to employ a Detail Mini-Nets block in the second stage, responsible for learning refinement displacement field that encodes high-frequency geometry details.
@@ 2.3 Fine-Grained Neural Representations for Geometry Modeling
Recent research [65, 67, 68] has brought attention to a persistent challenge in neural networks that the learned models tend to favor the retention of low-frequency signals while neglecting high-frequency information, known as the spectral bias [65]. In the context of geometry modeling, recent advancements have introduced hybrid approaches that combine implicit and explicit 3D representations, e.g. , voxels [69, 70, 71], tri-plane [72, 73] and hash-grids [74, 75, 76], to mitigate the issue above. In particular, Voxurf [71] employed voxel-based representations combined with an implicit function defined by a shallow MLP to represent a scene. It also introduced hierarchical geometry features to enhance perception across voxels. EG3D [72] presented a memory-efficient tri-plane architecture that scales efficiently with resolution, yielding higher capacity for handling complex 3D data. TensoRF [77] adopted a unique approach to decompose the 3D volume into vector-matrix outer products, leading to improved rendering quality. Additionally, INGP [74] combined voxel grids and multi-resolution hash tables to create compact and adaptive data representations. This innovative approach resulted in significant enhancements in both rendering quality and processing speed. However, it is worth noting that while these hybrid 3D representations offer advantages in various applications, using them as 3DMMs can be challenging due to issues related to the discontinuity of the underlying explicit 3D representations. Such discontinuity may result in instability and incorrect training outcomes, particularly when modeling complex surfaces such as facial shapes.
@@
Another emerging research trend seeks to bolster the intrinsic capabilities of neural networks for improved details. Inspired by the natural language processing field, NeRF [49] utilized position encoding to render finer-grained details. In addition, SIREN [78] and GARF [79] introduced innovative network activation functions, utilizing Sin and Gaussian functions, respectively. However, training these networks can be a challenging endeavor due to instability and susceptibility to over-fitting. To overcome this, FreeNeRF [80] and Nerfies [81] proposed to use a coarse-to-fine learning strategy. Furthermore, IDF [82] and HF-NeuS [83] extended the conventional explicit and discrete displacement mapping into ${\mathbb{R}}^{3}$ domain to capture high-frequency details within single static scenes. Nevertheless, there is relatively limited exploration into the modeling of dynamic high-frequency facial details across different individuals. To enhance the ability of INRs to build 3DMMs, our method employs a decomposition of the learning process. This approach begins by automatically learning deformations linked to identity and expression while simultaneously establishing correspondences between different individuals. It then learns to encode dynamic high-frequency details within the established template space, significantly facilitating the representation capability.

### v3

@@ Abstract
Accurate representations of 3D faces are of paramount importance in various computer vision and graphics applications. However, the challenges persist due to the limitations imposed by data discretization and model linearity, which hinder the precise capture of identity and expression clues in current studies. This paper presents a novel 3D morphable face model, named ImFace++, to learn a sophisticated and continuous space with implicit neural representations. ImFace++ first constructs two explicitly disentangled deformation fields to model complex shapes associated with identities and expressions, respectively, which simultaneously facilitate automatic learning of point-to-point correspondences across diverse facial shapes. To capture more sophisticated facial details, a refinement displacement field within the template space is further incorporated, enabling fine-grained learning of individual-specific facial details. Furthermore, a Neural Blend-Field is designed to reinforce the representation capabilities through adaptive blending of an array of local fields. In addition to ImFace++, we devise an improved learning strategy to extend expression embeddings, allowing for a broader range of expression variations. Comprehensive qualitative and quantitative evaluation demonstrates that ImFace++ significantly advances the state-of-the-art in terms of both face reconstruction fidelity and correspondence accuracy.
@@ 1 Introduction
In 3DMMs, the most fundamental issue lies in the way to generate latent morphable representations, and during the past two decades, along with data improvement in scale, diversity, and quality [8, 11, 6, 41], remarkable progresses have been achieved. The methods are initially linear model based [1, 57, 58] and further multilinear model based [9, 7, 56], where different modes are individually encoded. Unfortunately, for the relatively limited representation ability of linear models, these methods are not so competent at handling the cases with complex variations, e.g. , exaggerated expressions. In the era of deep learning, a number of nonlinear models have been investigated with the input of 2D images [12, 19] or 3D meshes [14, 15, 16, 17, 44] by using Convolutional Neural Networks (CNN) or Graph Neural Networks (GNN). They indeed deliver performance gains; however, restricted by the resolution of discrete representing strategies on input data, facial priors are not sufficiently captured, incurring loss of shape details. Besides, all current methods are dependent on the preposed procedure of point-to-point correspondence [10, 59, 60, 43], but face registration itself remains challenging.
@@
Figure. 1: ImFace++ encodes intricate variations associated with identities and expressions by two explicitly disentangled deformation fields with respect to a template face. Leveraging the established point-to-point correspondence, it applies an additional refinement displacement field to capture high-frequency details in the individual-independent template space. The model finally yields sophisticated and implicit representations that are highly adaptable for 3D face modeling.
@@
Recently, several studies have illuminated the potential of Implicit Neural Representations (INRs) [20, 21, 22, 26, 27] in precise modeling of 3D geometries. INRs describe input observations as low-dimensional shape embeddings and estimate the Signed Distance Function (SDF) or the occupancy values of query points, allowing for the definition of surfaces with arbitrary resolutions and topologies through isocontours. Such continuous representations offer advantages over the traditional discrete ones like voxels, point clouds, and meshes, and have reported substantial successes in shape reconstruction [62, 61, 28, 63] as well as surface registration [23, 25, 24]. This property suggests an alternative to 3DMM that can fulfill accurate shape correspondence and geometry modeling in a unified network. Despite that, applying INRs to face modeling presents unique challenges. Unlike the objects with distinct shapes and limited non-rigid variations, such as indoor scenes and human bodies, facial surfaces are inherently similar but also encompass intricate deformations arising from the interweaving of multiple identities and rich expressions. Current INR methods struggle to precisely capture and disentangle the intricacies, as evidenced by the preliminary attempt  [33].
@@
On the other side, it is essential to emphasize the faithful reproduction of facial details. In the context of geometry modeling, there has been a growing interest in hybrid approaches that combine explicit and implicit representations to harness the strengths of both techniques [74, 72, 69], aiming to address the spectral bias [65] issue often observed in neural networks and learn more sophisticated geometry clues. However, incorporating explicit 3D representations into INRs can introduce discontinuities, potentially leading to training instability and over-fitting [91, 103], especially when applied to complex structures like facial surfaces. An alternative research trend focuses on incorporating computer graphics techniques, such as displacement mapping [82, 83], into INRs to encode rich details. While these methods have shown remarkable performance in reconstructing single static scenes, the challenge remains in developing INR models capable of effectively capturing dynamically changing high-frequency details across multiple facial shapes, such as wrinkles and dimples.
@@
In response to the aforementioned challenges, this paper presents a novel two-stage 3D face morphable model, namely ImFace++, which substantially upgrades conventional 3DMMs by integrating customized INRs. In its initial stage, ImFace++ tackles nonlinear facial deformation modeling and establishment of point-to-point correspondence among different facial shapes by creating separate INR sub-networks, which explicitly disentangle shape morphs into two deformation fields for identity and expression, respectively (as Fig. 1 shows). The template shape is defined by a deep SDF, therefore eliminating the need for the face registration phase required by numerical solutions. To build more sophisticated 3DMMs, ImFace++ further incorporates a Refinement Displacement Field (RDF) in the second stage to capture high-frequency geometry details present in various faces. Notably, this stage for detail modeling is specifically tailored to adapt to the dynamic nature of facial expressions. By leveraging the established correspondences, RDF is learned within the individual-independent template space characterized by its relative stability, allowing the model to concentrate on encoding subtle details. Overall, by learning the disentangled deformation fields and RDF, ImFace++ accurately represents inter-individual changes and fine-grained clues while retaining the flexibility for its application in related tasks.
@@
Inspired by linear blend skinning [37], we further introduce a Neural Blend-Field to decompose the entire facial deformation or geometry into semantically meaningful regions encoded by an array of local implicit functions and adaptively blend them through a lightweight module, leading to more advanced representations with reduced parameter overheads. In addition to ImFace++, an improved embedding learning strategy is devised to extend the latent space of expressions to allow a more diverse range.
Overall, this study makes the following contributions:
@@
  • A novel INR-based 3DMM, namely ImFace++. It encodes complex facial shape variations with identity and expression through explicitly disentangled deformation fields, facilitating learning flexible and semantically meaningful representations.
@@
  • A refinement displacement field, i.e. RDF, within the individual-independent template space. Leveraging the established correspondence, the field faithfully encodes high-frequency facial details, thus enhancing the representation capabilities of ImFace++.
@@
  • A Neural Blend-Field. It enables fine-grained learning by adaptively blending an array of local implicit functions.
@@
  • Compelling results in different tasks. Extensive experiments evidence that ImFace++ excels in synthesizing high-quality 3D faces with intricate details, outperforming the state-of-the-art alternatives.
@@
A preliminary version of this study, i.e. ImFace, previously published in [87]. This paper significantly improves it in the following aspects: (i) We extend the model to a two-stage framework. To sufficiently capture facial priors, the upgraded model leverages the correspondences between different individuals, achieved by employing disentangled deformation fields in the first stage as in ImFace [87], and further applies a newly designed refinement field to progressively model face geometry, significantly alleviating the loss of shape details while relieving the burden of model learning. (ii) We update the model to reconstruct and synthesize 3D faces spanning a wide range of facial expressions and it renders more vivid results. We also refine the evaluation in both the observed and latent representation spaces, along with more comprehensive discussions. (iii) To further validate the proposed approach, we conduct comparison with other INR-based methods [33, 66] in terms of 3D face reconstruction and face correspondence. (iv) We showcase applications in various scenarios, including facial expression editing, detail transfer, multi-view reconstruction, and ear-to-ear face reconstruction, to demonstrate the adaptability of ImFace++ in the realm of 3D face modeling.
@@
The rest of this paper is organized as follows. Sec. 2 reviews related work on 3DMMs and INRs. Sec. 3 elaborates on the proposed ImFace++ method. Sec. 4 presents and analyzes the experimental results, followed by Sec. 5, which concludes this paper and provides insights for future perspectives.
@@ 2 Related Work
The published studies closely related to our study can be primarily summarized into: (i) 3D Morphable Models (3DMMs), (ii) Implicit Neural Representations (INRs), in particular those aiming to capture complex 3D structures, and (iii) fine-grained neural representations for geometry modeling. In the following, we provide a concise review of the relevant literature.
@@
The concept of 3DMMs was first introduced by Blanz and Vetter [1] as a foundation framework for face representation. This approach involves registering a known template mesh to all training scans using the Non-rigid Iterative Closest Point (NICP) algorithm [40]. Principal Components Analysis (PCA) is subsequently employed to create prior face distributions. To account for identity-dependent expressions, 3DMMs were extended to multilinear models [9, 7, 56]. Over time, great progress has been made aided by improvements in data quality [8, 11, 6, 41]. Notably, the Facial LAndmarks and More Expressions (FLAME) model [6] was introduced as an expressive model, capable of controlling facial expressions through a combination of jaw articulation and linear expression blendshapes. Trained on a comprehensive 3D facial dataset, including D3DFACS [42], FLAME delivers remarkable results, although it faces challenges in accurately capturing nonlinear facial deformations.
Deep neural networks have significantly accelerated the development of more powerful nonlinear 3DMMs. A number of models have been trained directly from 2D images [12, 19, 18, 101], but they often struggle to capture high-fidelity and fine-grained details. Such a limitation is attributed to the inherent challenges posed by the low resolution of input images in solving this ill-posed inverse problem. Recognizing the need to better leverage 3D face scans, [14] devises a method for mapping 3D meshes to the 2D space. Subsequently, several studies [44, 15, 17, 16] pursue the learning of 3DMMs from meshes, employing spectral or spiral convolutions. ASM [102] presents an adaptive skinning model, which utilizes the Gaussian Mixture Model (GMM) to generate skinning weights. While promising, these neural networks are established on discrete 3D representations, which tend to fail when confronting with intricate facial deformations and high-frequency details.
@@ 2.2 Implicit Neural Representations
In recent years, implicit neural functions have emerged as powerful and effective representations for 3D geometries [20, 21, 22, 26, 27, 45]. These approaches revolutionize the field by enabling continuous and highly expressive representations of geometric structures without the need for discretization. To address the complexity of shape variations and correspondence relationships, some attempts further incorporate an additional implicit deformation latent space [23, 24].
A number of studies focus on applying INRs to modeling human heads or bodies [35, 34, 33, 32, 31, 36, 30, 47, 66, 88, 89, 90, 92, 93, 95, 94]. In particular, NPMs [95] and GNPM [94] employ two neural networks to separately model the body shape and pose spaces. Models such as H3D-Net [34], SIDER [100], and SIRA++[92] learn implicit priors for 3D head reconstruction. i3DMM [33] makes significant strides by pioneering a generative implicit 3DMM tailored for human heads. However, utilizing neural networks to generate fine-grained details from raw signals often results in over-smoothed or implausible facial shapes. A concurrent work, NPHM [66], introduces a parametric head model using local implicit fields, and it adopts local INRs for identity representation and a globally conditioned INR for modeling expression deformations. Nonetheless, its ability for capturing high-frequency facial details is constrained by the globally conditioned neural network architecture. Another series of works [88, 89, 90, 93] concentrate on modeling facial appearances, leveraging techniques like differential volume rendering [49] and auto-decoding [20] to learn morphable radiance fields. But these models represent the underlying geometry using density, leading to noisy geometry approximations.
Figure. 2: ImFace++ overview. The model is constructed through two-stage learning. (a) The first stage primarily focuses on nonlinear deformation modeling and surface correspondence between different faces. The network comprises three Mini-Nets blocks designed to explicitly disentangle shape morphs into separate deformation fields, where the Expression and Identity Mini-Nets blocks are associated with expression and identity deformations, respectively. Meanwhile, the Template Mini-Nets block learns the SDF of a template face shape. (b) With the point-to-point correspondence established, ImFace++ proceeds to employ a Detail Mini-Nets block in the second stage, responsible for learning refinement displacement field that encodes high-frequency geometry details.
@@ 2.3 Fine-Grained Neural Representations for Geometry Modeling
Recent research [65, 67, 68] brings attention to a persistent challenge in neural networks that the learned models tend to favor the retention of low-frequency signals while neglecting high-frequency information, known as the spectral bias [65]. In the context of geometry modeling, latest advancements introduce hybrid approaches that combine implicit and explicit 3D representations, e.g. voxels [69, 70, 71], tri-planes [72, 73] and hash-grids [74, 75, 76], to mitigate the issue above. In particular, Voxurf [71] employs voxel-based representations combined with an implicit function defined by a shallow MLP to represent a scene. It also builds hierarchical geometry features to enhance perception across voxels. EG3D [72] presents a memory-efficient tri-plane architecture that efficiently scales with resolution, yielding a higher capacity for handling complex 3D data. TensoRF [77] decomposes the 3D volume into vector-matrix outer products, leading to improved rendering quality. Additionally, INGP [74] combines voxel grids and multi-resolution hash tables to create compact and adaptive data representations. This approach results in substantial enhancements in both rendering quality and processing speed. Inspired by the domain of signal processing, NeuRBF [99] and NFFB [98] utilize Radial Basis Functions (RBF) and filter banks as irregular grids to improve the capability. However, using these hybrid 3D representations as 3DMMs is challenging due to discontinuities in explicit 3D representations, incurring inaccurate gradients and poor generalizability, as discussed in the supplementary material of iNGP [74] and [103].
@@
Another research trend seeks to bolster the intrinsic capabilities of neural networks for improved details. Inspired by the natural language processing field, NeRF [49] applies position encoding to render finer-grained details. In addition, SIREN [78], GARF [79], and WIRE [97] introduce various network activation functions, including Sin, Gaussian, and Gabor wavelet functions. However, training these networks is not so straightforward due to instability and susceptibility to over-fitting. To overcome this, FreeNeRF [80] and Nerfies [81] add a coarse-to-fine learning strategy. Furthermore, IDF [82] and HF-NeuS [83] extend the conventional explicit and discrete displacement mapping into the ${\mathbb{R}}^{3}$ domain to capture high-frequency details within single static scenes. Nevertheless, there is relatively limited exploration into the modeling of dynamic high-frequency facial details across different individuals. To enhance the ability of INRs for 3DMMs, our method employs a decomposition of the learning process, which begins by automatically learning deformations linked to identity and expression while simultaneously establishing correspondences between different individuals. It then learns to encode dynamic high-frequency details in the established template space, significantly facilitating the representation capability.